OpenStudy (anonymous):

solve by factoring 6x^2 - 5x=50

ganeshie8 (ganeshie8):

what have you tried so far ?

OpenStudy (haichi):

11x^2

OpenStudy (midhun.madhu1987):

Try using the Quadratic Formula...

OpenStudy (mathmath333):

\(\large\tt \color{black}{ax^2+bx+c=(x-\alpha)(x-\beta)}\) \(\large\tt \color{black}{where~~\alpha=\dfrac{-b+\sqrt{b^2-4ac}}{2a}}\) \(\large\tt \color{black}{where~~\beta=\dfrac{-b-\sqrt{b^2-4ac}}{2a}}\)

OpenStudy (midhun.madhu1987):

This is another method :  take common factors from the first bracket and the second one.